16. Jennie and Alex both wanted to get a free ticket for a College Music concert. However, the concert staff told them the tickets were limited. Twenty people wanted to attend the concert but only 10 free tickets were left. So the concert center staff decided to use a lottery to decide who would receive the free tickets. Whats the probability of Jennie and Alex both getting free tickets? (5 pts)
Answer #1

Since there are 10 open positions and 20 candidates, so the probability of getting selected = 10/20 = 0.5

Each person has the same probability of getting the free ticket, equal to 0.5

Total possible ways of choosing 10 people out of 20 = 20C10

Now, assume that both of them get selected for free tickets.

The number of possible combinations for this to happen = 2C2*18C8

So the required probability = Number of possible ways for the case of interest/Total possible ways of selection = (2C2*18C8)/(20C10) = 9/38 = 0.237